OpenStudy (anonymous):

Prove the following, preferably using induction so i can learn it: \[\sum_{i=1}^{n}{2i-1}=n^2\]

OpenStudy (amistre64):

induction, step 1, prove its true for some solid value

OpenStudy (anonymous):

provwe ir's true for n = 1 :-D

OpenStudy (amistre64):

yep; 2(1)-1 = 1 is true

OpenStudy (amistre64):

next step, assume its true for an arbitrary value k

OpenStudy (anonymous):

yeah! the book also uses the letter k!

myininaya (myininaya):

then show it is true for k+1

OpenStudy (anonymous):

hmm how do i do that part?

OpenStudy (amistre64):

\[\sum_{1}^{k}2k-1=k^2\]

OpenStudy (amistre64):

then add 2(k+1)-1 to both sides

OpenStudy (zarkon):

@amistre64 be careful with your notation

OpenStudy (anonymous):

how do i prove that the statement is true for n=k?

OpenStudy (amistre64):

will try ;)

OpenStudy (amistre64):

you assume that k is any value; and you know its good for k=1

OpenStudy (anonymous):

so i just assume it's good for k, and then try to prove for n=k+1?

OpenStudy (amistre64):

yep

myininaya (myininaya):

\[k^2+2k+1=(k+1)^2\]

OpenStudy (amistre64):

if you can see a spot that it is bad for, then it a given that this wont work

OpenStudy (amistre64):

but there is no restriction to the domain as is other than >=1

myininaya (myininaya):

\[\sum_{i=1}^{k+1}(2i-1)=\sum_{i=1}^{k}(2i-1)+2(k+1)-1=k^2+2(k+1)-1 =(k+1)^2\]

myininaya (myininaya):

any questions agdgdgdwngo? ok what does rn and dn mean in computer programming?

OpenStudy (anonymous):

I still don't get step 2

myininaya (myininaya):

assume it is true for some k?

OpenStudy (amistre64):

proof by induction is a more rigid version of the "and so on and so forth" of days gone by

myininaya (myininaya):

step 1 show it is true for first k step 2 assume it is true for some k step 3 use step 2 to show it is true for k+1

OpenStudy (anonymous):

oh alright I get it... and after step 2 you basically prove the statement is true for an infinite number of cases

OpenStudy (amistre64):

yep, because if its true for any arbitrary k, and also true for the next one after it; that covers them all

OpenStudy (anonymous):

this type of proof can be easily done by a computer program.

myininaya (myininaya):

would this program require rn and dn?

OpenStudy (jamesj):

btw, it's a good idea to write down exactly the statement you're trying to prove, as it will make it very clear in your argument what it is you're doing. In this problem, I would write the statement as following. Let P(n) be the statement that \[ \sum_{i=1}^n (2i - 1) = n^2 \] Now. A proof by induction of this or any other such statement has two steps, as has been mentioned. Formalizing this idea, the steps are: 1. Show that P(1) is true. (Or P(m) for some other integer m) 2. Show that P(k) => P(k+1) for all \( k \geq 1 \) (Or \( k \geq m \) ) If you show these two things, then by the Principle of Mathematical Induction P(n) is true for all \( n \geq 1 \) (or \( n \geq m \) ).

OpenStudy (anonymous):

probably not :-P never heard of rn and dn.

myininaya (myininaya):

:(

myininaya (myininaya):

google doesn't work

OpenStudy (jamesj):

Now in this case, step 1 is indeed easy. P(1) is the statement that \[ \sum_{i=1}^1 (2i-1) = 1 \] and this is true.

myininaya (myininaya):

i know up means press u at first i thought it meant up

OpenStudy (jamesj):

In step 2, we start with the hypothesis P(k). And now you need to show that P(k) implies P(k+1). The algebra for showing that has already been written down. =========

myininaya (myininaya):

that was my algebra i can do a little algebra :)

OpenStudy (anonymous):

hmm... I think I get it.

myininaya (myininaya):

Prove \[\sum_{i=1}^{n}i=\frac{n(n+1)}{2}\]

OpenStudy (anonymous):

hmm alright... so first we prove P(1) by showing that 1 = 1

OpenStudy (anonymous):

and then we prove that if P(2),...,P(k) is true, then P(k+1) is also true

myininaya (myininaya):

yes

OpenStudy (anonymous):

so alright, let's prove P(2), which is basically saying that 1 + 2 = 0.5 * 6

OpenStudy (jamesj):

The way to write step 2 formally would be something like this: Let us suppose P(k). I.e., \[ \sum_{i=1}^k (2i - 1) = k^2 \] Now we want to show this implies P(k+1). Now \[ \sum_{i=1}^{k+1} (2i - 1) = \sum_{i=1}^{k} (2i - 1) + \sum_{i=k+1}^{k+1} (2i - 1) \] \[ = k^2 + \sum_{i=k+1}^{k+1} (2i - 1) \ \ \hbox{, by hypothesis P(k) } \] \[ = k^2 + (2k + 1) \] \[ = (k+1)^2 \] Now as k was arbitrary, P(k) => P(k+1).

myininaya (myininaya):

you only need to do it for one n=1 or n=2 but you don't need both and then assume it is true for k and now show it is true for k+! \[\sum_{i=1}^{k+1}i=\sum_{i=1}^{k}i+[(k+1)] \]

myininaya (myininaya):

but what does \[\sum_{i=1}^{k}i =? \]

OpenStudy (anonymous):

it is 1 + 2 + ... + k

myininaya (myininaya):

which is what did we get to assume is true: this right: \[\sum_{i=1}^{k}i=\frac{k(k+1)}{2}\] ?

OpenStudy (anonymous):

I'm confused.

OpenStudy (anonymous):

let's just show that \[\sum_{i=1}^{k+1}i = \sum_{i=1}^{k}{i} + k + 1\]

OpenStudy (anonymous):

or am I doing it backwards?

myininaya (myininaya):

\[step \text{ 1 : } P(1)=1\] \[step: \text{2 : } P(k)=\frac{k(k+1)}{2}\] \[step: 3 : \text{ use step 2 \to show} \text{ (in this case:)}\] \[ P(k+1)=P(k)+k+1=\frac{(k+1)([k+1]+1)}{2}\]

OpenStudy (anonymous):

I think we substitute that sum of i with \[\frac{k(k+1)}{2}\] and then we show

OpenStudy (jamesj):

So, in notation. P(n) is the statement \[ \sum_{i=1}^n i = n(n+1)/2 \] Now suppose P(k) i.e., \[ \sum_{i=1}^k i = k(k+1)/2 \] Then \[ \sum_{i=1}^{k+1} i =\sum_{i=1}^k i + (k+1) \] \[ = k(k+1)/2 + (k+1), \ \ \ \hbox{ by P(k) } \] \[ = \frac{ k^2 + k + 2k + 2}{2} \] \[ = \frac{(k+1)(k+2)}{2} \] I.e., P(k) => P(k+1) for arbitrary k. Therefore, as 1. P(1) is true 2. P(k) => P(k+1) for arbitrary \( k \geq 1 \) it follows by the PMI that P(n) is true for all \( n \geq 1 \).

OpenStudy (anonymous):

alright I understand proof by mathematical induction!

OpenStudy (anonymous):

first we set k to 1 and then prove P(k)

OpenStudy (jamesj):

No.

OpenStudy (anonymous):

no? :(

OpenStudy (jamesj):

Go back and read my statement of the method carefully.

OpenStudy (anonymous):

Proof by mathematical induction: _________________________________________________________ 1. Show that P(1) is true. (Or P(m) for some other integer m) 2. Show that P(k) => P(k+1) for all k≥1 (Or k≥m )

OpenStudy (jamesj):

Yes.

OpenStudy (anonymous):

I get it now :-D

OpenStudy (jamesj):

Like just about everything in mathematics, you'll know you really get it when you've written out quite a few problems.

OpenStudy (anonymous):

alright I have another problem coming up